OpenStudy (anonymous):

Consider F(x)= "integral from x to 2" (1/(1+t^4)). Find F'(x).

OpenStudy (anonymous):

\[F(x)=\int_x^2\frac{1}{1+t^4}dt\]

OpenStudy (anonymous):

Sorry, sorry, i meant to say from 2 to x but I just need a general idea of how to do this kind of problem so continue.

OpenStudy (anonymous):

\[F'(x)=\frac{d}{dt}\int_x^2\frac{1}{1+t^4}dt=\frac{d}{dt}(F(2)-F(x))\]\]

OpenStudy (anonymous):

\[\frac{d}{dt}(F(x)-F(2))\]

OpenStudy (anonymous):

so in other words you're going to get F(2) which derivative is zero

OpenStudy (anonymous):

now you have to worry about F(x) and use the chain rule

OpenStudy (anonymous):

\[F(x)=F(x)-F(2)\] \[F'(x)=f(x)\frac{d}{dx}x=1f(x)=\frac{1}{1+x^4}\]